OpenStudy (anonymous):

The following pictures depict objects whose sizes have different orders of magnitude. The approximate sizes are indicated in meters and kilometers. The sizes in the above sequence, converted to meters, make a geometric sequence. What is the ratio of the sequence? A. 1 B. 0.1 C. 0.01 D. 10 E. 100

OpenStudy (anonymous):

@jenny1994

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (butterflydreamer):

the question is basically asking you to find the ratio of the geometric series in metres :) So looking at the picture, we have the following sequence: 1, 10, 100 ,1000, 10 000, 100 000, 1000 000, 10 000 000 To find the ratio all you need to do is : 100/10 = 10/1 = ?

OpenStudy (anonymous):

d???

OpenStudy (butterflydreamer):

yeep :)

OpenStudy (anonymous):

thx

OpenStudy (butterflydreamer):

you're welcome ^_^